# An In-depth Look at Gemini’s Language Abilities

### 18 Dec 2023 v1

@@ 6.1 Experimental Details
We follow the standard zero-shot code evaluation pipeline provided by the ODEX [^8] https://github.com/zorazrw/odex. We take its recommended hyperparameters with a temperature of 0.8 and top_p of 0.95. We also add a customized instruction “Complete the given code with no more explanation. Remember that there is a 4-space indent before the first line of your generated code.” to ensure that the models’ output fits the desired format.
@@
We perform evaluation-based execution, measuring the Pass@1 metric, which determines whether a single sample from the model passes test cases (Chen et al., 2021). Since code generation is evaluated in a zero-shot fashion, the model may inevitably output code that does not conform to our input format well. Therefore, we perform rudimentary post-processing to regulate the output code and make it fit into the final verification pipeline as much as possible, including the removal of markdown code blocks, the extraction of function implementations and the truncation of stop tokens. It’s noteworthy that for incorrectly formatted indent, we do not manage to manually fix such issues and instead regard them as typical syntax errors.
@@ 6.2 Results and Analysis
In this section, we examine the overall performance and present a case study on the code generation.
First, from the overall results shown in Figure 15, we can see that Gemini Pro achieves a Pass@1 lower than GPT 3.5 Turbo and much lower than GPT 4 Turbo on both tasks. The results demonstrate that Gemini’s code generation capabilities still have room for improvement.
@@
We also present the analysis of how the libraries required in each solution affect the model performance in 15(b). Gemini Pro performs worse than GPT 3.5 on most library-used cases, such as mock, pandas, numpy, and datetime. However, it outperforms GPT 3.5 and GPT 4 on the matplotlib cases, showing stronger capabilities when performing drawing visualization via code.
@@
>>> remove_duplicates([1, 2, 3, 2, 4])
[1, 3, 4]
%**** 06_Code.tex Line 100 ****
@@ 7.1 Experimental Details
We investigate the efficacy of zero-shot and five-shot prompting strategies across 20 language pairs. Following the guidelines proposed by Gao et al. (2023b), we utilized designated prompts for both zero-shot and few-shot machine translation (MT), as exemplified in Table 8. These prompt settings have shown distinct advantages for users of large language models (LLMs), as noted by Robinson et al. (2023). Our experimental setup employed a top_p value of 1, a temperature of 0.3, a context_length of -1, and max_tokens 500 to optimize the performance of the translation model.
@@
To evaluate the outputs, we utilized the following metrics:
spBLEU: BLEU, a standard in machine translation evaluation, was employed (Papineni et al., 2002). We computed spBLEU scores following the methodology outlined in Goyal et al. (2022), using the sacreBLEU toolkit (Post, 2018) with the SPM-200 tokenizer (NLLB Team et al., 2022).
chrF2++: Our primary metric, chrF2++, leverages the implementation provided by sacreBLEU (Post, 2018). This choice is motivated by its ability to address some of BLEU’s limitations. For simplicity, we refer to this metric as chrF in our discussion (Popović, 2017).
@@ 7.2 Results and Analysis
Lang. Gemini Pro GPT 3.5 Turbo GPT 4 Turbo Google NLLB
ssw_Latin 5.71 11.15 37.20 - 43.30
sna_Latin 4.36 16.07 42.82 44.40 43.40
ckb_Arab 0.01 19.42 39.58 47.70 47.20
mag_Deva 40.37 38.91 23.83 - 58.50
ibo_Latin 5.26 14.05 41.57 43.50 41.40
hau_Latin 37.32 23.62 49.92 53.20 53.50
pbt_Arab 0.33 20.32 33.22 - 39.40
tam_Tamil 0.01 34.52 48.87 55.80 53.70
kat_Geor 0.04 34.21 43.21 51.40 48.10
gle_Latin 4.60 47.66 56.81 60.10 58.00
kmr_Latin 10.78 25.98 31.96 40.00 39.30
war_Latin 59.14 49.94 54.92 - 57.40
ajp_Arab 49.79 46.93 46.45 - 51.30
lim_Latin 39.90 40.30 41.32 - 47.90
ukr_Cryl 57.92 54.67 57.09 58.60 56.30
fra_Latin 71.46 71.15 70.92 72.70 69.70
lvs_Latin 59.67 55.01 58.05 - 54.80
ron_Latin 65.58 64.20 64.58 65.00 61.30
tpi_Latin 31.13 37.33 48.36 - 41.60
acm_Arab 48.47 44.50 40.71 - 31.90
Table 4: Machine translation performance (chRF (%) scores) across models for all languages using 0-shot prompt. Best scores are bolded, second best underlined.
Lang. Gemini Pro GPT 3.5 Turbo GPT 4 Turbo Google NLLB
ssw_Latin 19.74 7.62 38.07 - 43.30
sna_Latin 4.36 15.84 42.95 44.40 43.40
ckb_Arab 0.01 24.56 40.71 47.70 47.20
mag_Deva 47.54 39.25 45.33 - 58.50
ibo_Latin 5.26 16.29 41.65 43.50 41.40
hau_Latin 5.32 24.22 50.11 53.20 53.50
pbt_Arab 0.33 21.35 34.11 - 39.40
tam_Tamil 0.01 34.86 48.69 55.80 53.70
kat_Geor 0.04 33.61 43.17 51.40 48.10
gle_Latin 4.60 47.30 57.25 60.10 58.00
kmr_Latin 4.91 26.10 32.76 40.00 39.30
war_Latin 48.94 50.85 56.26 - 57.40
ajp_Arab 50.72 47.49 48.12 - 51.30
lim_Latin 46.92 43.25 45.21 - 47.90
ukr_Cryl 57.79 55.19 56.85 58.60 56.30
fra_Latin 71.80 71.34 70.79 72.70 69.70
lvs_Latin 59.93 55.05 58.34 - 54.80
ron_Latin 66.11 64.19 64.42 65.00 61.30
tpi_Latin 35.90 37.39 50.67 - 41.60
acm_Arab 49.78 45.88 46.45 - 31.90
Table 5: Machine translation performance (chRF (%) scores) models for all languages using 5-shot prompt. Best scores are bolded, second best underlined.
@@
In Table 4 and Table 5, we conduct a comparative analysis of Gemini Pro, GPT 3.5 Turbo, and GPT 4 Turbo against established systems like Google Translate. [^9] http://translate.google.com Additionally, we benchmark against NLLB-MoE (NLLB Team et al., 2022), the leading open-source machine translation (MT) model known for its extensive language coverage. The results indicate that Google Translate generally outperforms other models, and excels in 9 languages, followed by NLLB which excels on 6 and 8 language in 0/5-shot settings. The general language models showed competitive performances but have not yet surpassed the dedicated machine translation systems in translation into non-English languages.
Figure 17: Machine translation performance (chRF (%) scores) by language pairs
Figure 17 illustrates the comparative performance of general language models across language pairs. GPT 4 Turbo showed a consistent deviation of performance with NLLB relative to GPT 3.5 Turbo and Gemini Pro. This reflects the findings in the literature on GPT 4 Turbo’s multilingual performance OpenAI (2023). GPT 4 Turbo also offered larger improvements for low-resource languages (as measured in NLLB Team et al. (2022)), whereas for high-resource languages performance was similar between the LLMs. In comparison, Gemini Pro outperforms both GPT 3.5 Turbo and GPT 4 Turbo on 8 out of 20 languages, and achieved the top performances on 4 languages. However, Gemini Pro showed a strong tendency to to block responses in approximately 10 language pairs, which we further study in the next analysis.
Figure 18: Number of samples that are blocked by Gemini Pro
Figure 19: Performance in chrf (%) on blocked and unblocked samples
Figure 20: Performance (chrf (%)) by script
@@
Figure 18 highlights that Gemini Pro’s lower performance in these languages is due to its tendency to block responses in scenarios of lower confidence. A response is deemed "blocked" if Gemini Pro in either its 0-shot or 5-shot configuration generates a Blocked Response error. A closer examination in Figure 19 reveals that Gemini Pro marginally outperforms GPT 3.5 Turbo and GPT 4 Turbo in unblocked samples where it demonstrates higher confidence. Specifically, it surpasses GPT 4 Turbo by 1.6 chrf in 5-shot and 2.6 chrf in 0-shot settings, and exceeds GPT 3.5 Turbo by 2.7 chrf and 2 chrf in 5-shot and 0-shot settings, respectively. However, our initial analysis of GPT 4 Turbo and GPT 3.5 Turbo’s performance on these samples indicates they are typically more challenging to translate. Gemini Pro’s subpar performance on these particular samples is especially apparent in instances where the Gemini Pro 0-shot blocks responses but the 5-shot does not, and vice versa.
@@
Throughout our analysis of the models, we observed that few-shot prompts generally yield a modest enhancement in average performance, with an increasing variance pattern following the order: GPT 4 Turbo < GPT 3.5 Turbo < Gemini Pro. While Gemini Pro’s 5-shot prompts show improvement over its 0-shot counterparts in languages where it demonstrates confidence, in certain languages, such as hau_Latin, the model exhibits significantly reduced confidence, resulting in blocked responses (refer to Table 5).
In Figure 20, we present apparent trends when categorizing languages by family or script. A key observation is Gemini Pro’s competitive performance with other models on Cyrillic scripts, is contrasted by its underperformance on other scripts. GPT-4 stands out, outperforming other models across various scripts, with few-shot prompts being particularly effective. This effectiveness is especially pronounced in languages using the Devanagari script.
Figure 21: Web agent success rate of evaluateioned models at different site groups
@@ 8 Web Agents   [Uncaptioned image]
Gemini-pro 7.09 3.52
Gemini-pro 5.23 4.83
GPT-3.5-turbo 8.75 6.44
GPT-3.5-turbo 6.41 6.06
GPT-4-turbo 15.16 14.22
@@
Finally, we examine the ability of each model to act as a web navigation agent, a task that requires long-term planning and complex data understanding. We use WebArena (Zhou et al., 2023b), an execution-based simulation environment where the success criterion is based on execution outcome. Tasks given to agents consist of information seeking, site navigation, and content & configuration operations. The tasks span over a variety of web sites, including E-commerce platforms, social forums, collaborative software development platforms (e.g. gitlab), content management systems, and online maps.
@@ 8.1 Experiment Details
In sum, we tested with two prompts from WebArena: p_cot_id_actree_2s and p_cot_id_actree_2s_no_na, which are respectively CoT prompt with the UA hint and CoT prompt without the UA hint. To make results comparable between GPTs and Gemini, we set the same upper limit on the observation lengths for all of them. This number is set to 1920 tokens using the tokenizer of gpt-4-1106-preview, consistent with experiments in WebArena. In terms of hyper-parameters, we used the default suggested by each of the large language model providers. For the Gemini models, the suggested default temperature is 0.9 and default top-p is 1.0, and the WebArena’s suggested default for GPT models is 1.0 for temperature and 0.9 for top-p.
@@
The action sequence of an agent is considered correct as long as they achieved the final goal, regardless the intermediate steps they take. We use WebArena’s evaluation, which determines wether a task is completed successfully or not with the agent’s final output. A small number of responses were blocked by the Gemini API (around 2% of the total test cases), and we treat these as failed trajectories in our experiments.
@@ 8.2 Results and Analysis
In general, Gemini-Pro predicts more tasks as unachievable, especially in the case where a UA hint is given, as shown in Figure 22. Gemini-Pro predicts over 80.6% of the tasks as unachievable when given an UA hint, compared to 47.7% by GPT-3.5-Turbo. Note that 4.4% of the tasks in the dataset are actually unachievable, so both far over-predict the actual number of unachievable tasks.
@@
At the same time, we observed that Gemini Pro has a higher tendency to respond in shorter phrases and take fewer steps before reaching a conclusion. As shown in 22(a), more than half of trajectories by Gemini Pro are under ten steps, while majority of trajectories by GPT 3.5 Turbo and GPT 4 Turbo are between 10 and 30 steps. Similarly, the majority of Gemini responses are less than 100 characters in length, while most of GPT 3.5 Turbo, GPT 4 Turbo, and Mixtral’s responses are over 300 characters in length 22(b). Gemini tends to directly predict the actions while other models would start with reasoning and then give their action predictions.
@@ 9 Conclusion
  • The Gemini Pro model, which is comparable to GPT 3.5 Turbo in model size and class, generally achieves accuracy that is comparable but somewhat inferior to GPT 3.5 Turbo, and much worse than GPT 4. It outperforms Mixtral on every task that we examined.
  • In particular, we find that Gemini Pro was somewhat less performant than GPT 3.5 Turbo on average, but in particular had issues of bias to response order in multiple-choice questions, mathematical reasoning with large digits, premature termination of agentive tasks, as well as failed responses due to aggressive content filtering.
  • On the other hand, there were bright points: Gemini performed better than GPT 3.5 Turbo on particularly long and complex reasoning tasks, and also was adept multilingually in tasks where responses were not filtered.
@@
First, our work is a snapshot in time with respect to ever-changing and unstable API-based systems. All results here are current as of this writing on December 18, 2023, but may change in the future as models and the surrounding systems are upgraded.
Second, the results may be dependent on the specific prompts and generation parameters that we selected. It is quite possible that with further prompt engineering, or multiple samples and self-consistency as was used by Gemini Team (2023), the results could change significantly. However, we believe that the consistent results over several tasks with standardized prompts is a reasonable indication of the robustness and generalized instruction following capability of the tested models.
@@
Based on this paper, we can make the recommendation to researchers and practitioners to carefully look at the Gemini Pro model as a tool in the toolbox, comparable to GPT 3.5 Turbo. Gemini’s Ultra edition, which is yet to be released, is reported to be on par with GPT 4, and a further examination of this model will be warranted when it is available.

### 24 Dec 2023 v2

@@ 6.1 Experimental Details
We follow the standard zero-shot code evaluation pipeline provided by the ODEX [^9] https://github.com/zorazrw/odex. We generate with temperature 0.0, which demonstrated the best performance for all models, compared to other temperatures. We use a prompt of “Complete the following python3 function: ” to ensure that the models’ output fits the desired format.
@@
We perform execution-based evaluation, measuring the Pass@1 metric, which determines whether a single model output passes test cases (Chen et al., 2021). Since code generation is evaluated in a zero-shot fashion, the model may inevitably output code that does not conform to our input format well. Therefore, we perform rudimentary post-processing to make the output code fit into the final verification pipeline as much as possible, including the removal of markdown code blocks, the extraction of function implementations and the truncation of natural language sentences following the code. We do not automatically fix missing library imports.
@@ 6.2 Results and Analysis
In this section, we examine the overall performance and present a few examples.
First, from the results shown in Figure 15, we can see that Gemini Pro achieves a Pass@1 lower than GPT 3.5 Turbo and GPT 4 Turbo on both tasks. On HumanEval, Gemini Pro significantly outperforms Mixtral, and on ODEX Mixtral and Gemini Pro have roughly equivalent accuracy. [^10] Note that the 59.8% accuracy that Gemini Pro achieved on HumanEval is significantly lower than that reported by Gemini Team (2023). We suspect the difference may be due to differences in prompting techniques, and will continue examining the issue.
@@
On the ODEX benchmark all models achieved lower accuracy than HumanEval, with a significant portion of the errors being due to the models failing to import libraries that they were using, or using non-current APIs. We also present the analysis of how the libraries required in each solution affect the model performance in 15(b). Gemini Pro performs slightly worse than GPT 3.5 on most library-used cases, such as mock, pandas, numpy, and datetime.
@@
$>$ remove_duplicates([1, 2, 3, 2, 4])
%**** 06_Code.tex Line 100 ****[1, 3, 4]
@@ 7.1 Experimental Details
We use a five-shot prompting strategy (zero shot results are also noted in Appendix C), specifically the prompts proposed by Gao et al. (2023b) designated in Table 7. Our experimental setup employed a top_p value of 1, a temperature of 0.3, a context_length of -1, and max_tokens 500, which we found to generally achieve good performance for translation.
@@
To evaluate the outputs, we utilized sentence level averaged chrF2++, leveraging the implementation provided by sacreBLEU (Post, 2018). This standard metric is based on character and word $n$-gram overlap between the system output and the reference sentence. We compute the sentence level chrF scores. For simplicity, we refer to this metric as chrF in our discussion (Popović, 2017).
@@ 7.2 Results and Analysis
Lang. Gemini Pro GPT 3.5 Turbo GPT 4 Turbo Mixtral Google NLLB
ssw_Latin ${}^{\dagger}$ 18.16 38.62 18.84 - 43.51
sna_Latin ${}^{\dagger}$ 24.44 43.25 22.20 44.36 43.40
ckb_Arab ${}^{\dagger}$ 26.69 41.30 18.36 47.61 47.25
mag_Deva 34.54 39.70 45.46 25.93 - 58.03
ibo_Latin ${}^{\dagger}$ 21.46 41.94 17.75 43.37 41.36
hau_Latin ${}^{\dagger}$ 30.24 50.82 22.47 53.18 53.25
pbt_Arab ${}^{\dagger}$ 22.81 34.21 16.61 - 39.27
tam_Taml ${}^{\dagger}$ 35.50 48.04 23.78 55.98 53.63
kat_Geor ${}^{\dagger}$ 33.32 40.94 23.78 51.11 47.10
gle_Latin ${}^{\dagger}$ 46.72 56.52 26.93 59.93 57.87
kmr_Latin ${}^{\dagger}$ 30.03 33.33 19.04 39.94 39.25
war_Latin ${}^{\dagger}$ 51.17 56.01 34.74 - 57.25
ajp_Arab 50.64 47.45 47.01 33.60 - 50.61
lim_Latin 39.99 43.77 46.05 32.31 - 47.58
ukr_Cyrl 56.89 54.56 56.44 49.66 58.25 56.04
fra_Latin 70.77 70.99 70.77 66.73 72.98 70.01
lvs_Latin 59.49 54.55 57.95 31.17 62.49 54.89
ron_Latin 65.09 63.18 63.93 56.68 65.08 61.21
tpi_Latin ${}^{\dagger}$ 40.14 47.97 $3$ 42.02
acm_Arab 49.05 45.26 44.44 31.65 - 32.60
Table 4: Machine translation performance (chrF (%) scores) across models for all languages using 5-shot prompt. Best scores are bolded, second best underlined. ${}^{\dagger}$ indicates languages where Gemini Pro blocked more than 50% of responses.
@@
In Table 4, we conduct a comparative analysis of Gemini Pro, GPT 3.5 Turbo, GPT 4 Turbo, and Mixtral. We also compare against established translation-specific systems like Google Translate [^11] http://translate.google.com, and NLLB-MoE (NLLB Team et al., 2022), the leading open-source machine translation (MT) model known for its extensive language coverage.
The results indicate that Google Translate generally outperforms other models in languages that it supports, excelling in 10 languages. It is followed by NLLB, which excels on 6 languages. Gemini Pro provided impressive accuracy on several languages, even having the best accuracy on 3: South Levantine Arabic, Romanian, and Mesopotamian Arabic (surpassing not only GPT 3.5 Turbo, but also GPT 4 Turbo).
However, on average, the results indicate that the general-purpose language models showed competitive performances but have not yet surpassed the dedicated machine translation systems in translation into non-English languages.
Figure 17: Machine translation performance (chrF (%) scores) by language pairs for 5-shot prompt.
Figure 17 illustrates the comparative performance of language models across language pairs for the 5-shot prompt. GPT 4 Turbo showed a consistent deviation of performance with NLLB relative to GPT 3.5 Turbo and Gemini Pro. This reflects the findings in the literature on GPT 4 Turbo’s multilingual performance (OpenAI, 2023). GPT 4 Turbo also offered larger improvements for low-resource languages (as measured by NLLB Team et al. (2022)), whereas for high-resource languages performance was similar between the LLMs. In comparison, Gemini Pro outperforms both GPT 3.5 Turbo and GPT 4 Turbo on 5 out of 20 languages, and achieved the top performances on 3 languages. Mixtral underperformed the other models across language pairs.
Figure 18: Number of samples that are blocked by Gemini Pro 5-shot
Figure 19: Performance in chrf (%) on blocked and unblocked languages for 5-shot prompt.
@@
If we analyze responses at a language level, we see in Figure 18 that Gemini Pro’s lower performance in 12/20 languages is due to its tendency to block responses on particular languages, generally ones with lower levels of resources. We consider a response "blocked" if Gemini Pro generates a Blocked Response error, and define unblocked languages as those languages where >50% samples are not blocked.
Examining performance at a language level in Figure 19, we see that Gemini Pro outperforms GPT 3.5 Turbo and GPT 4 Turbo on 5/8 unblocked languages most of which are high resource. Additionally, in Figure Figure 19, we observe that the implementation of safety filters leads to a decrease in the overall chrF score. This reduction occurs because the filters block samples from languages that the model otherwise handles relatively effectively.
(a) chrf by target sentence length
(b) chrf by predicted sentence length
Figure 20: Performance with varying target and predicted length using 5-shot prompt on unblocked languages.
@@
In Figure 21, we present apparent trends when categorizing languages by family or script. A key observation is Gemini Pro’s competitive performance with other models on Cyrillic scripts, is contrasted by its underperformance on other scripts. GPT 4 Turbo stands out, outperforming other models across various scripts, especially in the Devanagari script.
Figure 21: Performance by script
In Figure 20, we examine the performance of various models across different sentence length segments, categorized by both target length (19(a)) and predicted length (19(b)). Upon scrutinizing Figure 20, we observe that Gemini Pro’s performance at longer target lengths does not match that of GPT 4 Turbo and GPT 3.5 Turbo. However, when considering predicted lengths, Gemini Pro generally outperforms both GPT 4 Turbo and GPT 3.5 Turbo at longer lengths, suggesting it produces higher quality translations at longer lengths. Additionally, the figure suggests that a significant portion of the performance decline at shorter predicted lengths even on unblocked languages may be attributed to empty predictions, likely triggered by the content filtering mechanism.
@@ 8 Web Agents   [Uncaptioned image]
Gemini-pro 7.12 3.52
Gemini-pro 6.25 4.83
GPT-3.5-turbo 8.87 6.44
GPT-3.5-turbo 6.36 6.06
GPT-4-turbo 14.90 14.22
@@
Finally, we examine the ability of each model to act as an instruction following agent that performs tasks on the web, which requires long-term planning and complex data understanding. We use WebArena (Zhou et al., 2023b), an execution-based simulation environment where the success criterion is based on execution outcome. Tasks given to agents consist of information seeking, site navigation, and content & configuration operations. The tasks span over a variety of web sites, including E-commerce platforms, social forums, collaborative software development platforms (e.g. gitlab), content management systems, and online maps.
@@ 8.1 Experiment Details
In sum, we tested with two prompts from WebArena: p_cot_id_actree_2s and p_cot_id_actree_2s_no_na, which are the CoT prompt with the UA hint and CoT prompt without the UA hint, respectively. To make results comparable between GPTs and Gemini, we set the same upper limit on the observation lengths for all of them. This number is set to 1920 tokens using the tokenizer of gpt-4-1106-preview, consistent with experiments in WebArena. In terms of hyper-parameters, we used the default suggested by each of the large language model providers. For the Gemini models, the suggested default temperature is 0.9 and default top-p is 1.0, and the WebArena suggested default for GPT models is 1.0 for temperature and 0.9 for top-p.
@@
The action sequence of an agent is considered correct as long as they achieved the final goal, regardless the intermediate steps they take. We use WebArena’s evaluation, which determines wether a task is completed successfully or not with the agent’s final output.
@@ 8.2 Results and Analysis
Figure 23: Web agent success rate of evaluated models at different site groups
In general, Gemini-Pro predicts more tasks as unachievable, especially in the case where a UA hint is given, as shown in Figure 22. Gemini-Pro predicts over 80.6% of the tasks as unachievable when given a UA hint, compared to 47.7% by GPT-3.5-Turbo. Note that 4.4% of the tasks in the dataset are actually unachievable, so both far over-predict the actual number of unachievable tasks.
@@
At the same time, we observed that Gemini Pro has a greater tendency to respond in shorter phrases and take fewer steps before reaching a conclusion. As shown in 23(a), more than half of trajectories by Gemini Pro are under ten steps, while majority of trajectories by GPT 3.5 Turbo and GPT 4 Turbo are between 10 and 30 steps. Similarly, the majority of Gemini responses are less than 100 characters in length, while most of GPT 3.5 Turbo, GPT 4 Turbo, and Mixtral’s responses are over 300 characters in length 23(b). Gemini tends to directly predict the actions while other models would start with reasoning and then give their action predictions.
@@ 9 Conclusion
  • The Gemini Pro model, which is comparable to GPT 3.5 Turbo in model size and class, generally achieves accuracy that is comparable but somewhat inferior to GPT 3.5 Turbo, and much worse than GPT 4 on English tasks.
  • In particular, we find that Gemini Pro was somewhat less performant than GPT 3.5 Turbo on average, but in particular had issues of bias to response order in multiple-choice questions, mathematical reasoning with large digits, and premature termination of agentive tasks. When using the default content filtering settings, there were also failed responses due to aggressive content filtering.
  • On the other hand, there were bright points: Gemini performed better than GPT 3.5 Turbo on particularly long and complex reasoning tasks.
  • In addition, when generating text in other languages (specifically through translation), Gemini Pro outperforms both GPT 3.5 Turbo and GPT 4 Turbo on the languages where requests are not blocked, but there are several languages for which Gemini Pro does not return any answer.
  • The open-source model Mixtral is competitive with Gemini Pro and GPT 3.5 Turbo on Knowledge-based QA and Mathematics tasks, but significantly underperformed on other tasks.
@@
First, our work is a snapshot in time with respect to ever-changing and unstable API-based systems. All results here are current as of this writing on December 24, 2023, but may change in the future as models and the surrounding systems are upgraded.
Second, the results may be dependent on the specific prompts and generation parameters that we selected. In fact, we found that the results of all models were affected by prompt selection, and that the GPT models seemed somewhat more robust to small variations in the prompts of the GPT models. It is quite possible that with further prompt engineering, or multiple samples and self-consistency as was used by Gemini Team (2023), the results could change significantly.
@@
Based on this paper, we can make the recommendation to researchers and practitioners to carefully look at the Gemini Pro model as a tool in the toolbox, comparable to GPT 3.5 Turbo. In particular, Gemini Pro may be a preferable alternative when processing non-English languages. Gemini’s Ultra edition, which is yet to be released, is reported to be on par with GPT 4, and a further examination of this model will be warranted when it is available.